OpenStudy (anonymous):

Question One: If Julie’s group decides that they need the limo service for 3 hours which is most economical company to use? We can solve for both equations using 3 for x. Remember from our word problem, x is the hours rented. Go-Limo Lime Limo Y = 300 + 100x Y = 300 + 100(3) Y = 300 + 300 Y = 600 Y = 200 + 150x Y = 200 + 150(3) Y = 200 + 450 Y = 650 For a three hour rental the Go-Limo would be the best buy. Question Two: If Julie’s wants to find the time where one company would be a better buy than the other she can use the substitution method. Go-Limo Lime Limo Y = (300 + 1

OpenStudy (anonymous):

y=300+100x y=200+150x substitution means that you express y or x from 1 of the equations and plug that in to the other. Lets use this: y=300+100x 300+100x=200+150x 100=50x x=2

OpenStudy (anonymous):

meaning that if the work takes 2 hours than the price is = for the 2 companies

OpenStudy (anonymous):

Hope this is what you needed :-)